OpenStudy (anonymous):

how do you get 20 out of writing 60% AS A FRACTION IN LOWEST TERMS

OpenStudy (radar):

60%=60/100=3/5 Hey you got me i don't see "20" Check that problem out again and see if you are overlooking something